4. Find the equation of the tangent line to curve 3 (x2 + y2) ? = 25 (x2 - y2),= 1 at (2,1).
Answer #1

we have

3(x^2+y^2)^2=25(x^2-y^2)

\frac{d}{dx}3(x^2+y^2)^2=\frac{d}{dx}25(x^2-y^2)

3\cdot 2(x^2+y^2)\left (2x+2y\frac{dy}{dx} \right )=25\left (2x-2y\frac{dy}{dx} \right )

12x(x^2+y^2)+12y(x^2+y^2)\frac{dy}{dx}=50x-50y\frac{dy}{dx}

12y(x^2+y^2)\frac{dy}{dx}+50y\frac{dy}{dx}=50x-12x(x^2+y^2)

(12y(x^2+y^2)+50y)\frac{dy}{dx}=50x-12x(x^2+y^2)

\frac{dy}{dx}=\frac{50x-12x(x^2+y^2)}{12y(x^2+y^2)+50y}

\frac{dy}{dx}=\frac{25x-6x(x^2+y^2)}{6y(x^2+y^2)+25y}

the slope of the tangent line is,

dy m de

m=\frac{25x-6x(x^2+y^2)}{6y(x^2+y^2)+25y}

slope at (2,1),

m=\frac{25(2)-6(2)(2^2+1^2)}{6(1)(2^2+1^2)+25(1)}

m=\frac{50-12(5)}{6(5)+25}

m=\frac{50-60}{30+25}

m=-\frac{10}{55}

m=-\frac{2}{11}

the equation of the tangent line is,

Or - 2) = 0h - h

put the x0 = 2, y0 = 1 and m = -2/11

y-1=-\frac{2}{11}(x-2)

y=-\frac{2}{11}x+\frac{4}{11}+1

y=-\frac{2}{11}x+\frac{15}{11}